# Flow Matching Meets PDEs: A Unified Framework for Physics-Constrained Generation

### v2

@@ 6 Conclusions
Additionally, our results highlight the benefits of stochastic sampling strategies, which outperform deterministic methods in cases involving complex target distributions. Beyond these specific benchmarks, the proposed framework generalizes to a wide class of PDE-constrained problems formulated via residual minimization. By combining the computational efficiency of flow matching with the interpretability and rigor of physics-based modeling, PBFM offers a powerful and flexible tool for surrogate modeling, uncertainty quantification, and simulation acceleration in physics and engineering applications.
@@ Appendix A Flow Matching
Appendix B Dataset generation and residual computation
@@ B.1 Darcy flow
with $r=10$ and $w=0.125$. The permeability field $K(\boldsymbol{x})$ is modeled as $K(\boldsymbol{x})=\exp(G(\boldsymbol{x}))$, where $G(\boldsymbol{x})$ is a Gaussian random field. The pressure field is generated by solving a least-squares problem based on a 64-term truncated Karhunen–Loève expansion, following the approach of Jacobsen et al. (2025).
@@
B.2 Kolmogorov flow
We generate two distinct datasets (training and validation) for the Kolmogorov flow problem with Reynolds numbers in the range [100, 500][100,\ 500], using a spatial resolution of $128\times 128$. The simulation is based on the vorticity–stream function formulation. The velocity field is obtained from the computed vorticity through the stream function, while the pressure field is derived by solving the pressure Poisson equation. We employ TorchFSM (Liu et al., 2025b) to perform GPU-accelerated flow simulations using the Fourier spectral method. The training dataset includes 32 different flow conditions sampled via a Halton sequence (Kocis and Whiten, 1997), while the validation dataset contains 16 conditions. For each condition, 256 simulations are conducted with slightly perturbed initial states. Simulations are run for 10 000 time steps to reach a statistically steady state, followed by data sampling every 4 000 time steps. With a time step size of $dt=1/\text{Re}$, this yields 1 024 snapshots per condition.
@@ B.2 Kolmogorov flow
B.3 Dynamic stall
The dynamic stall datasets used for training and validation are generated by solving the unsteady, compressible, two-dimensional RANS equations around a sinusoidally pitching NACA0012 airfoil. An O-grid mesh is utilized, featuring 512 nodes along the airfoil surface and 128 nodes in the normal direction. Simulations are performed using ANSYS Fluent 2024R2 (ANSYS, 2024). The governing equations are discretized using a second-order upwind scheme for spatial accuracy and a second-order implicit scheme for time integration. Gradient reconstruction employs a least-squares cell-based method, while fluxes are computed using the Rhie-Chow momentum interpolation. Pressure-velocity coupling is handled through a coupled solver. Airfoil pitching is modeled by prescribing a rigid-body motion to the entire mesh, defined by the angular velocity function $\dot{\alpha}(t)=\omega\ \alpha_{s}\cos(\omega t)$. The mean angle of attack, $\alpha_{0}$, is applied at the start of each simulation by rotating the mesh accordingly. To close the RANS equations, the SST turbulence model with an intermittency transport equation is employed. Each oscillation cycle is resolved using 2 048 time steps, and simulations are run until periodic convergence is achieved.
@@ B.3 Dynamic stall
where $\mathcal{N}(0,0.02)$ denotes a Gaussian noise term with zero mean and standard deviation 0.02. This results in 32 perturbed variations per nominal condition, yielding a total of $128\times 32$ simulations for training and $16\times 32$ for testing. To reduce computational costs, all simulation fields are downsampled to a resolution of $128\times 128$. The saved quantities include fields of absolute pressure, density, temperature, signed skin friction, and tangential velocity gradients across the airfoil surface over a full pitching cycle. Figure 9 illustrates a spatio-temporal representation of the wall shear stress ($\tau_{w}$) across one complete pitching cycle, highlighting the pitch-up and pitch-down phases and demonstrating the structured mapping of surface data over time.
@@
Appendix C Additional samples and analysis
We conclude the analysis of the proposed test cases by presenting additional comparisons and complete field prediction examples. We compute the 1D Wasserstein-1 distance to evaluate distributional differences. This approach treats each pixel independently rather than modeling the full 2D spatial structure.
C.1 Darcy flow
Figure 10 shows the pressure and permeability fields for the different methods under analysis. To ensure a fair comparison of the residuals, all fields have similar variable ranges, since higher magnitudes can lead to disproportionately large residual values. This choice is further justified by the residual plots, which exhibit higher absolute errors in regions with large permeability. Despite local differences, a clear trend emerges: starting from the FM baseline, where residual peaks reach values around 100, to the PBFM method, where peak residuals are reduced by approximately an order of magnitude.
@@ C.1 Darcy flow
To conclude our analysis of the Darcy flow, Figure 11 compares the DiT architecture with the UNet used by Bastek et al. (2025). This comparison corresponds to the FM setup without residual loss. In the first panel, which reports the FM loss during training, we observe that the UNet architecture reproduces the overfitting behavior noted by Bastek et al. (2025). In contrast, the DiT architecture exhibits a stable and monotonic reduction in training loss. The second panel shows the residual error, which remains comparable between the two models overall, with the exception of slight discrepancies at one function evaluation. Finally, in the third panel, the Wasserstein distance reveals a clear gap in performance: the UNet consistently incurs higher errors, with the distance being approximately twice as large for pressure and up to four times larger for permeability, underscoring the superior distributional accuracy of the DiT model. Additional samples produced by our method are shown in Figure 12.
@@
C.2 Kolmogorov flow
Figure 13 illustrates example predictions from the FM, PBFM-1, and PBFM models, along with their corresponding physical residuals, representing the divergence of the predicted fields. The baseline FM exhibits a residual MSE on the order of $10^{-1}$ across most of the domain, whereas the unrolled framework significantly reduces the error, dropping below $10^{-3}$ in large regions of the field, despite some localized areas with higher residuals.
@@ C.2 Kolmogorov flow
Furthermore, Figure 14 shows the mean and standard deviation of the predicted fields for all models, alongside the reference data. All frameworks closely match the reference mean, with the unrolled version providing a smoother, less oscillatory solution compared to FM and PBFM-1. Regarding the standard deviation, none of the models fully capture the reference distribution, particularly missing the peak value of approximately $0.95$, although their performance remains broadly comparable.
@@
Additional samples produced by PBFM method are shown in Figure 15.
@@
C.3 Dynamic stall
For the dynamic stall case, Figure 16 presents the mean and standard deviation of the predicted fields obtained using PBFM. The predictions show strong agreement with the reference data, both in terms of overall distribution as well as capturing the extrema of each variable. The only notable deviation appears in the standard deviation of the skin friction, which slightly exceeds the corresponding reference values.
@@ C.3 Dynamic stall
Figure 17 compares the performance of the proposed models as a function of the number of FM integration steps. The first panel focuses on the physical residual MSE, indicating that the best performance is achieved with 10 FM steps, with PBFM consistently delivering an $8\times$ reduction compared to the baseline. The second panel reports the Wasserstein distance, showing that PBFM also narrows the gap with the reference data distribution, reaching its optimal value at 20 FM steps. Finally, the MSE of the mean and standard deviation reveals that the proposed methods preserve the already strong performance of the baseline, introducing only minor deviations.
Additional qualitative samples produced by our trained PBFM method are shown in Figure 18. They highlight the wide range of physical behavior modeled by this case: from small oscillation attached flow at the top, to deep dynamic stall cases with shock wave formation at the bottom.
@@
Appendix D Residual loss scaling laws
@@
Figure 19 presents a comparison of the MSE for physical residuals, as well as the mean and standard deviation, for both the PBFM-1 and PBFM frameworks. The results show that unrolling helps regularize the error, producing a monotonic increase in error as a function of the power $p$, and also reduces sensitivity to the choice of $p$ in the range $[1,4]$. Notably, both frameworks achieve optimal performance when residuals are scaled linearly with time, $p=1$. In contrast, using no scaling, $p=0$, results in significantly higher errors, underscoring the importance of appropriately weighting residuals based on the starting time.
@@
Appendix E Conflict-free updates and weighted loss terms
@@
We evaluate the proposed setup on the dynamic stall case, the most challenging scenario considered, comparing ConFIG to models trained with various fixed loss weights. Figure 20 reports the MSE for the physical residual, the predicted mean, and the standard deviation across all configurations. ConFIG consistently outperforms the fixed-weight approaches, achieving the lowest error in each individual metric and delivering the best overall performance.
@@
Appendix F Architecture and training details
@@
We adopt the Diffusion Transformer (DiT) architecture proposed by Peebles and Xie (2023) as the backbone for our flow matching model, incorporating minor modifications. The model is conditioned via adaptive layer normalization (adaLN-Zero) blocks, which replace the standard normalization layers. The scale and shift parameters in these blocks are derived from the sum of the embedding vectors for the time step $t$, used in the flow matching process, and the conditioning signal $c$. We introduce two modifications compared to the original DiT implementation. First, we incorporate linear attention (Wang et al., 2020) alongside the standard quadratic one. Second, we replace the original label embedder with a two-layer module: a linear layer followed by a SiLU activation function, and a second linear layer that produces the final conditioning embedding.
@@
Appendix G Performance evaluation
We provide an overview of the training performance, focusing on the impact of the residual loss along with ConFIG, as well as the effect of unrolling. Table 3 reports the wall-clock time, in seconds, for a single training iteration on an NVIDIA A100 64 GB GPU. Incorporating ConFIG (Liu et al., 2025a) increases the training time by approximately 80% compared to the baseline FM model, primarily due to the additional backward pass required to compute gradients associated with the residual loss. Similarly, unrolling negatively affects training performance: using four unroll steps roughly doubles the iteration time. Importantly, while training becomes more expensive, inference time remains unchanged, an essential requirement for deploying fast and reliable surrogate models.
Table 3: Comparison of wall-clock time in seconds for one training iteration on an NVIDIA A100 64GB GPU for the proposed approaches. Batch size is 64 for all cases. Inference time is unchanged.
FM PBFM-1 PBFM PBFM PBFM
2 steps 3 steps 4 steps
Darcy flow $4.29\cdot 10^{-2}$ $8.14\cdot 10^{-2}$ $1.18\cdot 10^{-1}$ $1.55\cdot 10^{-1}$ $1.90\cdot 10^{-1}$
Kolmogorov flow $1.13\cdot 10^{-1}$ $1.94\cdot 10^{-1}$ $3.02\cdot 10^{-1}$ $4.10\cdot 10^{-1}$ $5.18\cdot 10^{-1}$
Dynamic stall $4.28\cdot 10^{-2}$ $6.69\cdot 10^{-2}$ $9.78\cdot 10^{-2}$ $1.28\cdot 10^{-1}$ $1.59\cdot 10^{-1}$
Table 4 reports the inference wall-clock times for generating 32, 64, and 128 dynamic stall samples using 8 CPU cores of an Intel Xeon Platinum 8358 and an NVIDIA A100 64GB GPU. The same CPU was used to run the numerical simulations, ensuring a more consistent comparison of computational performance. On average, a single numerical simulation requires 76 minutes to complete. In contrast, the proposed generative model can produce 128 samples in just 2 and 4 minutes when using 10 and 20 FM steps on the CPU, respectively. When executed on a modern GPU, these inference times drop dramatically to 0.2 and 0.4 seconds, respectively. This substantial speedup highlights the model’s potential as a fast and reliable surrogate for dynamic stall prediction in helicopter and wind turbine applications.
Table 4: Wall-clock time in seconds to generate $n$ samples using 8 cores of an Intel Xeon Platinum 8358 and an NVIDIA A100 64GB for dynamic stall case. Batch size is set equal to the number of samples. One simulation with the same cores takes on average 76 minutes ($4.56\cdot 10^{3}$ s).

### v3

@@ Discussion
In this paper, we introduced PBFM, a generative framework designed to improve physical consistency while preserving the strengths of the flow matching approach for high-dimensional data generation. Our model provides straightforward improvements over standard FM to minimize physical residuals, arising from PDEs or algebraic constraints, in a conflict-free manner. We conduct extensive benchmarks across three representative physical systems, demonstrating the versatility and robustness of our approach. The incorporation of temporal unrolling plays a key role, enabling improved final state approximations and mitigating Jensen’s gap by providing a more accurate prediction of the final sample. Additionally, our results highlight the benefits of stochastic sampling strategies, which outperform deterministic methods in cases involving complex target distributions. Beyond these specific benchmarks, the proposed framework generalizes to a wide class of PDE-constrained problems formulated via residual minimization. By combining the computational efficiency of flow matching with the interpretability and rigor of physics-based modeling, PBFM offers a powerful and flexible tool for surrogate modeling, uncertainty quantification, and simulation acceleration in physics and engineering applications.
@@ Appendix A Flow Matching
Appendix B Ablation with a toy problem
To provide an intuition for the proposed improvements, we consider an ablation with a toy problem where the neural network outputs are the $x$ and $y$ coordinates of points on a circumference subject to the physical constraint of unit radius. Fig. 7 shows the resulting point distribution along with the corresponding absolute physical error through mean estimation of the final sample. We compare with diffusion models DM, as a DDPM representative (Song et al., 2022), and PIDM-ME, the best performing physics-informed algorithm (Bastek et al., 2025), and a variant of the flow matching approach that uses Config without unrolling. It is apparent that each extension yields a noticeable gain, and the final PBFM model exhibits a residual error that is 61.8 and 27.5 times lower than the DM and PIDM-ME baselines.
Appendix C Dataset generation and residual computation
@@ C.1 Darcy flow
with $r=10$ and $w=0.125$. The permeability field $K(\boldsymbol{x})$ is modeled as $K(\boldsymbol{x})=\exp(G(\boldsymbol{x}))$, where $G(\boldsymbol{x})$ is a Gaussian random field. The pressure field is generated by solving a least-squares problem based on a 64-term truncated Karhunen-Loève expansion, following the approach of Jacobsen et al. (2025).
@@
C.2 Kolmogorov flow
We generate two distinct datasets (training and validation) for the Kolmogorov flow problem with Reynolds numbers in the range [100, 500][100,\ 500], using a spatial resolution of $128\times 128$. The simulation is based on the vorticity–stream function formulation. The velocity field is obtained from the computed vorticity through the stream function, while the pressure field is derived by solving the pressure Poisson equation. We employ TorchFSM (Liu et al., 2025b) to perform GPU-accelerated flow simulations using the Fourier spectral method. The training dataset includes 32 different flow conditions sampled via a Halton sequence (Kocis & Whiten, 1997), while the validation dataset contains 16 conditions. For each condition, 256 simulations are conducted with slightly perturbed initial states. Simulations are run for 10 000 time steps to reach a statistically steady state, followed by data sampling every 4 000 time steps. With a time step size of $dt=1/\text{Re}$, this yields 1 024 snapshots per condition.
@@ C.2 Kolmogorov flow
C.3 Dynamic stall
The dynamic stall datasets used for training and validation are generated by solving the unsteady, compressible, two-dimensional RANS equations around a sinusoidally pitching NACA0012 airfoil. An O-grid mesh is utilized, featuring 512 nodes along the airfoil surface and 128 nodes in the normal direction. Simulations are performed using ANSYS Fluent 2024R2 (ANSYS, 2024). The governing equations are discretized using a second-order upwind scheme for spatial accuracy and a second-order implicit scheme for time integration. Gradient reconstruction employs a least-squares cell-based method, while fluxes are computed using the Rhie-Chow momentum interpolation. Pressure-velocity coupling is handled through a coupled solver. Airfoil pitching is modeled by prescribing a rigid-body motion to the entire mesh, defined by the angular velocity function $\dot{\alpha}(t)=\omega\ \alpha_{s}\cos(\omega t)$. The mean angle of attack, $\alpha_{0}$, is applied at the start of each simulation by rotating the mesh accordingly. To close the RANS equations, the SST turbulence model with an intermittency transport equation is employed. Each oscillation cycle is resolved using 2 048 time steps, and simulations are run until periodic convergence is achieved (Baldan & Guardone, 2024, 2025).
@@ C.3 Dynamic stall
where $\mathcal{N}(0,0.02)$ denotes a Gaussian noise term with zero mean and standard deviation 0.02. This results in 32 perturbed variations per nominal condition, yielding a total of $128\times 32$ simulations for training and $16\times 32$ for testing. To reduce computational costs, all simulation fields are downsampled to a resolution of $128\times 128$. The saved quantities include fields of absolute pressure, density, temperature, signed skin friction, and tangential velocity gradients across the airfoil surface over a full pitching cycle. Fig. 8 illustrates a spatio-temporal representation of the wall shear stress ($\tau_{w}$) across one complete pitching cycle, highlighting the pitch-up and pitch-down phases and demonstrating the structured mapping of surface data over time.
@@
Appendix D Additional samples and analysis
We conclude the analysis of the proposed test cases by presenting additional comparisons and complete field prediction examples.
D.1 Darcy flow
Fig. 9 shows the pressure and permeability fields for the FM-OT and PBFM methods under analysis. To ensure a fair comparison of the residuals, all fields have similar variable ranges, since higher magnitudes can lead to disproportionately large residual values. This choice is further justified by the residual plots, which exhibit higher absolute errors in regions with large permeability. Despite local differences, a clear difference emerges: the FM baseline presents residual peak values around 100, while PBFM method peak residuals are reduced by approximately an order of magnitude.
@@ D.1 Darcy flow
To conclude our analysis of the Darcy flow, Fig. 10 compares the DiT architecture with the UNet used by Bastek et al. (2025). This comparison corresponds to the FM setup without residual loss. In the first panel, which reports the FM-OT loss during training, we observe that the UNet architecture reproduces the overfitting behavior noted by Bastek et al. (2025). In contrast, the DiT architecture exhibits a stable and monotonic reduction in training loss. The second panel shows the residual error, which remains comparable between the two models overall, with the exception of slight discrepancies at one function evaluation. Finally, in the third panel, the Wasserstein distance reveals a clear gap in performance: the UNet consistently incurs higher errors, with the distance being approximately twice as large for pressure and up to four times larger for permeability, underscoring the superior distributional accuracy of the DiT model. Additional samples produced by our method are shown in Fig. 11.
@@
D.1.1 Non-dimensional scaling
Even though the role of scaling is well recognized in machine learning, its impact on physical residuals and distributional performance remains an important aspect to examine. We study this effect in the context of the Darcy flow problem by comparing two scaling strategies. In the first approach, each variable is normalized to have zero mean, yielding the following scaling values: $p_{\mathrm{mean}}=0.0,\ p_{\mathrm{std}}=0.576,\quad k_{\mathrm{mean}}=1.386,\ k_{\mathrm{std}}=10.64.$ The corresponding results are reported in Table 2. The second approach, denoted as the Scale version, rescales the pressure $p$ to the range [1,1][-1,1] and the permeability $k$ to $[0,1]$. The results for this setting are shown in Table 8. Across all methods, this alternative scaling consistently degrades performance, leading to higher physical residual errors and poorer distributional metrics.
Table 8: Generative performance metrics for Darcy flow problem over 1024 samples. RE: physical Residual MSE, WD: Wassserstein Distance, JS: Jensen-Shannon divergence, NFE: Number of function evaluations, IT: Inference wall-clock time. Best and second best results.
Dataset Metric PBFM Scale FM-OT Scale DiffusionPDE Scale D-Flow Scale
Darcy RE 2.260 5.551 4.307 2.417
WD $\cdot 10^{2}$ 0.185 0.076 0.075 0.437
JS $\cdot 10^{1}$ 0.218 0.152 0.142 0.353
NFE 20 20 20 20
IT $[s]$ 0.101 0.100 0.590 3.126
‡ D-Flow method is unstable, samples are filtered using RE <5<5 condition resulting in 403 valid ones.
D.2 Kolmogorov flow
Fig. 12 illustrates example predictions from the FM-OT, and PBFM models, along with their corresponding physical residuals, representing the divergence of the predicted fields. The baseline FM exhibits a residual MSE on the order of $10^{-1}$ across most of the domain, whereas the unrolled framework significantly reduces the error, dropping below $10^{-3}$ in large regions of the field, despite some localized areas with higher residuals.
@@ D.2 Kolmogorov flow
Furthermore, Fig. 13 shows the mean and standard deviation of the predicted fields for FM-OT and PBFM models, alongside the reference data. The frameworks closely match the reference mean, with the unrolled version providing a smoother, less oscillatory solution compared to FM-OT. Regarding the standard deviation, none of the models fully capture the reference distribution, particularly missing the peak value of approximately $0.95$, although their performance remains broadly comparable.
@@
Additional samples produced by PBFM method are shown in Fig. 14.
@@
D.3 Dynamic stall
For the dynamic stall case, Fig. 15 presents the mean and standard deviation of the predicted fields obtained using PBFM. The predictions show strong agreement with the reference data, both in terms of overall distribution as well as capturing the extrema of each variable. The only notable deviation appears in the standard deviation of the skin friction, which slightly exceeds the corresponding reference values.
@@ D.3 Dynamic stall
Fig. 16 compares the performance of the proposed models as a function of the number of FM integration steps. The first panel focuses on the physical residual MSE, indicating that the best performance is achieved with 10 FM steps, with PBFM consistently delivering a one order of magnitude reduction compared to the baseline. The second panel reports the Wasserstein distance, showing that PBFM also narrows the gap with the reference data distribution, reaching its optimal value at 20 FM steps. Finally, the MSE of the mean and standard deviation reveals that the proposed methods preserve the already strong performance of the baseline, introducing only minor deviations.
Additional qualitative samples produced by our trained PBFM method are shown in Fig. 17. They highlight the wide range of physical behavior modeled by this case: from small oscillation attached flow at the top, to deep dynamic stall cases with shock wave formation at the bottom.
@@
Appendix E Residual loss scaling laws
@@
Fig. 18 presents a comparison of the MSE for physical residuals, as well as the mean and standard deviation, for both the FM-OT with Config and PBFM frameworks. The results show that unrolling helps regularize the error, producing a monotonic increase in error as a function of the power $p$, and also reduces sensitivity to the choice of $p$ in the range $[1,4]$. Notably, both frameworks achieve optimal performance when residuals are scaled linearly with time, $p=1$. In contrast, using no scaling, $p=0$, results in significantly higher errors, underscoring the importance of appropriately weighting residuals based on the starting time.
@@
Appendix F Conflict-free updates and weighted loss terms
@@
We evaluate the proposed setup on the dynamic stall case, the most challenging scenario considered, comparing ConFIG to models trained with various fixed loss weights. Fig. 19 reports the MSE for the physical residual, the predicted mean, and the standard deviation across all configurations. ConFIG consistently outperforms the fixed-weight approaches, achieving the lowest error in each individual metric and delivering the best overall performance.
To further assess the effectiveness of ConFIG, we quantified gradient conflicts via pairwise cosine similarity between physical and FM losses. For the best model using 4-step unrolling with ConFIG, the average conflict is 6.93%, whereas a model with fixed residual scaling at 500 exhibits a much higher average conflict of 19.98%.
@@
Appendix G Architecture and training details
@@
We adopt the Diffusion Transformer (DiT) architecture proposed by Peebles & Xie (2023) as the backbone for our flow matching model, incorporating minor modifications. The model is conditioned via adaptive layer normalization (adaLN-Zero) blocks, which replace the standard normalization layers. The scale and shift parameters in these blocks are derived from the sum of the embedding vectors for the time step $t$, used in the flow matching process, and the conditioning signal $c$. We introduce two modifications compared to the original DiT implementation. First, we incorporate linear attention (Wang et al., 2020) alongside the standard quadratic one. Second, we replace the original label embedder with a two-layer module: a linear layer followed by a SiLU activation function, and a second linear layer that produces the final conditioning embedding.
@@
Appendix H Performance evaluation
Table 10 reports the inference wall-clock times for generating 32, 64, and 128 dynamic stall samples using 8 CPU cores of an Intel Xeon Platinum and an NVIDIA A100 64GB GPU. The same CPU was used to run the numerical simulations, ensuring a more consistent comparison of computational performance. On average, a single numerical simulation requires 76 minutes to complete. In contrast, the proposed generative model can produce 128 samples in just 2 and 4 minutes when using 10 and 20 FM steps on the CPU, respectively. When executed on a modern GPU, these inference times drop dramatically to 0.2 and 0.4 seconds, respectively. This substantial speedup highlights the model’s potential as a fast and reliable surrogate for dynamic stall prediction in helicopter and wind turbine applications.
Table 10: Wall-clock time in seconds to generate $n$ samples using 8 cores of an Intel Xeon Platinum and an NVIDIA A100 64GB for dynamic stall case. Batch size is set equal to the number of samples. One simulation with the same cores takes on average 76 minutes ($4.56\cdot 10^{3}$ s).